A person heading to work travels an average of 38 mph. it takes hime 15 mins to get there. How far did he travel? how do u figure this out?
Answer by Fombitz(32388)

Rate x Time = Distance
You have to make sure you use consistent units.
Rate is speed, 38 mph (distance in miles, time in hours).
Time is 15 minutes. Convert to hours. 60 minutes in 1 hour.
T=15/60=0.25 hours.
R*T=D
38 miles/hr*0.25 hrs=9.5 miles
.
.
.
You could also have converted the rate to miles per minute.
R=38 mile/hr*(1 hr/60 min)=0.6333 miles/min
Then
D=R*T=0.63333 miles/min*15 min=9.5 miles